OpenStudy (anonymous):

HELP 1 MEDAL What are the solutions to the following system of equations? y = x2 + 3x - 7 3x - y = -2

OpenStudy (anonymous):

replace the y in the second equation with "x2+3x-7" and that'll find you y. After that, substitute the answer into one of the equations and solve for x

OpenStudy (haseeb96):

apply same method here which i told you before

OpenStudy (haseeb96):

3x+2=y now put it in eq(1) 3x+2=x^2 +3x -7 x^2 +3x -3x -7 -2 =0 x^2 -9 = 0 x^2 = 9 taking square root here on both sides \[x=\pm3\] now first put x= 3 and x =-3 y =3x +2 y = 3(3)+2 or y = 3(-3)+2 y= 9+2 or y = -9+2 y=11 or y =-7 Hence the solution set is that ={ ( 3,11 ) , ( -3,-7 )}

OpenStudy (haseeb96):

marrylin tthis is the solurtion of it

OpenStudy (anonymous):

thank you so much!!!!